# The extension of Schwarzschild line element to include uniformly accelerated mass

### v3

@@ B.1 Inverse Rindler transformation
$ct$ $=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right),\hskip 28.45274ptx=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}$ (B.1)
@@
$ct=\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\frac{c^{2}}{a_{x}}\Bigg{\{}1+\frac{a_{x}}{c^{2}}\,\bar{x}\Bigg{\}}$ (B.2)
and $\frac{c^{2}}{a_{x}}\left(1+\frac{a_{x}}{c^{2}}\,x\right)=\frac{c^{2}}{a_{x}}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg{\{}1+\frac{a_{x}}{c^{2}}\,\bar{x}\Bigg{\}}$ (B.3)
Now dividing equation (B.2) by (B.3) we obtain:
$\frac{ct}{\frac{c^{2}}{a_{x}}\left(1+\frac{a_{x}}{c^{2}}\,x\right)}=\tanh\left(\frac{a_{x}\,t}{c}\right)\text{\,\,\,\,\,\,or }\,\,\,\,c\bar{t}=\frac{c^{2}}{a_{x}}\tanh^{-1}\left[\frac{ct}{\frac{c^{2}}{a_{x}}+x}\right]$ (B.4)
Again from equations (B.2) and (B.3) we can write as:
$\Bigg{[}\frac{c^{2}}{a_{x}}\left(1+\frac{a_{x}}{c^{2}}\,x\right)\Bigg{]}^{2}-c^{2}t^{2}$ $=\left(\frac{c^{2}}{a_{x}}\right)^{2}\Bigg{\{}1+\frac{a_{x}}{c^{2}}\,\bar{x}\Bigg{\}}^{2}\Bigg{[}\cosh^{2}\left(\frac{a_{x}\,\bar{t}}{c}\right)-\sinh^{2}\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg{]}$
@@
$\text{or\,\,\,\,\,\,}\bar{x}$ $=\sqrt{\Bigg{[}\left(\frac{c^{2}}{a_{x}}+x\right)\Bigg{]}^{2}-c^{2}t^{2}}-\frac{c^{2}}{a_{x}}$ (B.5)
Equations (B.4) and (B.5) represents the inverse Rindler transformation between an accelerated frame $\bar{K}$ and a rest frame $K$ along the $X$-direction.
B.2 Inverse Rindler transformation for acceleration in arbitrary direction
@@
$\vec{r}_{\parallel}$ $=\frac{\left(\vec{r}\cdot\vec{a}\right)}{|\vec{a}|^{2}}\vec{a}\,\,\hskip 28.45274pt\text{and\,\,\,\,\,\,\,}\vec{r}_{\perp}=\vec{r}-\frac{\left(\vec{r}\cdot\vec{a}\right)}{|\vec{a}|^{2}}\vec{a}$ (B.7)
@@
$\vec{\bar{r}}_{\parallel}$ $=\Bigg{[}\sqrt{\left(\frac{\left(\vec{r}\cdot\vec{a}\right)}{|\vec{a}|^{2}}+\frac{c^{2}}{|\vec{a}|^{2}}\right)^{2}-\frac{c^{2}t^{2}}{|\vec{a}|^{2}}}-\frac{c^{2}}{|\vec{a}|^{2}}\Bigg{]}\vec{a}$ (B.9)
@@
$=\frac{c^{2}}{a}\tanh^{-1}\Bigg{\{}\frac{a\,t}{c\Big{(}\frac{\vec{a}}{c^{2}}\cdot\vec{r}+1\Big{)}}\Bigg{\}}$ (B.11)
and the spatial coordinate transformation can be obtained from (B.9) and (B.10) (by putting $|\vec{a}|=a$):
@@
$=\vec{r}-\Bigg{[}\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}-\Bigg{\{}\left(\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}+\frac{c^{2}}{a}\right)^{2}-c^{2}t^{2}\Bigg{\}}^{1/2}-\frac{c^{2}}{a}\Bigg{]}\frac{\vec{a}}{a}$ (B.12)
The equations (B.11) and (B.12) are the Rindler transformation of coordinates in arbitrary direction.
Now we consider that $\vec{a_{x}}$, $\vec{a_{y}}$ and $\vec{a_{z}}$ respectively are the $X$, $Y$ and $Z$ components of the acceleration $\vec{a}$ such that $a=\sqrt{a_{x}^{2}+a_{y}^{2}+a_{z}^{2}}$. Then we obtain from eqn. (B.11) as below:
$c\bar{t}$ $=\frac{c^{2}}{a}\tanh^{-1}\Bigg{\{}\frac{a\,t}{c\Big{(}\frac{a_{x}}{c^{2}}\,x+\frac{a_{y}}{c^{2}}\,y+\frac{a_{z}}{c^{2}}\,z+1\Big{)}}\Bigg{\}}$ (B.13)
and from (B.12):
$\vec{\bar{x}}+\vec{\bar{y}}+\vec{\bar{z}}$ $=\vec{x}+\vec{y}+\vec{z}-\Bigg{[}\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}-\Bigg{\{}\left(\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}+\frac{c^{2}}{a}\right)^{2}-c^{2}t^{2}\Bigg{\}}^{1/2}-\frac{c^{2}}{a}\Bigg{]}\frac{\vec{a}}{a}$
@@
$-\frac{c^{2}t^{2}}{a^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg{]}\vec{a}_{z}$ (B.14)
@@
$\hskip 85.35826pt+\Bigg{\{}\left(\frac{1}{a^{2}}\left(x\,a_{x}+y\,a_{y}+z\,a_{z}\right)+\frac{c^{2}}{a^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg{]}a_{x}$ (B.15)
@@
$\hskip 85.35826pt+\Bigg{\{}\left(\frac{1}{a^{2}}\left(x\,a_{x}+y\,a_{y}+z\,a_{z}\right)+\frac{c^{2}}{a^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg{]}a_{y}$ (B.16)
@@
$\hskip 85.35826pt+\Bigg{\{}\left(\frac{1}{a^{2}}\left(x\,a_{x}+y\,a_{y}+z\,a_{z}\right)+\frac{c^{2}}{a^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg{]}a_{z}$ (B.17)
@@
$c\bar{t}$ $=\frac{c^{2}}{a_{xy}}\tanh^{-1}\Bigg{\{}\frac{a_{xy}\,t}{c\Big{(}\frac{a_{x}}{c^{2}}\,x+\frac{a_{y}}{c^{2}}\,y+1\Big{)}}\Bigg{\}}$ (B.18)
$\bar{x}$ $=x+\Bigg{[}-\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\Bigg{\{}\left(\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg{]}a_{x}$ (B.19)
$\bar{y}$ $=y+\Bigg{[}-\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\Bigg{\{}\left(\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg{]}a_{y}$ (B.20)
@@
$c\bar{t}$ $=\frac{c^{2}}{a_{x}}\tanh^{-1}\Bigg{\{}\frac{c\,t}{\Big{(}\frac{c^{2}}{a_{x}}+x\Big{)}}\Bigg{\}}$ (B.22)
$\bar{x}$ $=\Bigg{\{}\left(x+\frac{c^{2}}{a_{x}}\right)^{2}-c^{2}t^{2}\Bigg{\}}^{1/2}-\frac{c^{2}}{a_{x}}$ (B.23)
$\bar{y}$ $=y$ (B.24)
$\text{and\,\,\,\,}\bar{z}$ $=z$ (B.25)
@@
B.3 Rindler transformation for the frames with non-zero initial velocity
@@
$ct$ $=\gamma\left(ct^{\prime}+\frac{v_{x}}{c}\cdot x^{\prime}\right),\hskip 28.45274ptx=\gamma\left(x^{\prime}+v_{x}\,t^{\prime}\right),\hskip 28.45274pty$ $=y^{\prime}\,\,,\hskip 28.45274ptz=z^{\prime}$ (B.26)
@@
$ct^{\prime}$ $=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\text{,\,\,\,\,\,\,\,}x^{\prime}=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}$ (B.27)
@@
Here it is to be noted that as stated above the accelerated frame $\bar{K}$ has zero initial velocity with respect to the frame $K^{\prime}$, however the frame $K^{\prime}$ itself is moving with uniform velocity $v_{x}$, therefore according to the observer in the rest frame $K$, the accelerated frame has non-zero initial velocity. The coordinate transformation relation between the inertial rest frame $K$ and the accelerated frame $\bar{K}$ can be obtained by substituting the expressions of the coordinates in equation (B.27) into equation (B.26) as follows:
$ct$ $=\gamma\Bigg{[}\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cdot\Bigg{\{}\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}\Bigg{\}}\Bigg{]}$ (B.28)
$x$ $=\gamma\Bigg{[}\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}+\frac{v_{x}}{c}\,\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg{]}$ (B.29)
$y$ $=\bar{y}$ (B.30)
 $\text{\,\,\,\,\,and\,\,\,\,\,\,}z$ $=\bar{z}$ (B.31)
@@
Equations (B.28) and (B.29) can be rewritten as:
$\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)$ $=\gamma\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\Bigg{[}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg{]}$ (B.32)
$\text{and\,\,\,\,\,\,}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)$ $=\gamma\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\Bigg{[}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg{]}$ (B.33)
Now dividing (B.32) by (B.33) we obtain:
$\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}$ $=\frac{\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)}{\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)}=\frac{\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}}{1+\frac{v_{x}}{c}\cdot\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)}$ (B.34)
@@
$\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}=\frac{\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\tanh\left\{\tanh^{-1}\left(\frac{v_{x}}{c}\right)\right\}}{1+\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)\tanh\left\{\tanh^{-1}\left(\frac{v_{x}}{c}\right)\right\}}=\tanh\Bigg{\{}\frac{a_{x}\,\bar{t}}{c}+\tanh^{-1}\left(\frac{v_{x}}{c}\right)\Bigg{\}}$
or $\frac{a_{x}\,\bar{t}}{c}=\tanh^{-1}\Bigg{\{}\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}\Bigg{\}}-\tanh^{-1}\left(\frac{v_{x}}{c}\right)$ (B.35)
@@
$c\bar{t}=\frac{c^{2}}{a_{x}}\tanh^{-1}\left\{\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)-\frac{v_{x}}{c}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)-\frac{v_{x}}{c}\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}\right\}$ (B.36)
Additionally, squaring both sides of equations (B.32) and (B.33) we can obtain as follows:
$\left(x+\gamma\frac{c^{2}}{a_{x}}\right)^{2}-\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)^{2}$ $=\gamma^{2}\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)^{2}\Bigg{[}\Bigg{\{}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg{\}}^{2}$
$-\Bigg{\{}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg{\}}^{2}\Bigg{]}$

### v4

@@ A.1 Inverse Rindler transformation
$ct$ $=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right),\hskip 14.22636ptx=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}$ (A.1)
@@
$ct=\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\frac{c^{2}}{a_{x}}\Bigg\{1+\frac{a_{x}}{c^{2}}\,\bar{x}\Bigg\}$ (A.2)
and $\frac{c^{2}}{a_{x}}\left(1+\frac{a_{x}}{c^{2}}\,x\right)=\frac{c^{2}}{a_{x}}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg\{1+\frac{a_{x}}{c^{2}}\,\bar{x}\Bigg\}$ (A.3)
Now, dividing equation (A.2) by (A.3) we obtain:
$\frac{ct}{\frac{c^{2}}{a_{x}}\left(1+\frac{a_{x}}{c^{2}}\,x\right)}=\tanh\left(\frac{a_{x}\,t}{c}\right)\text{\,\,\,\,\,\,or }\,\,\,\,c\bar{t}=\frac{c^{2}}{a_{x}}\tanh^{-1}\left[\frac{ct}{\frac{c^{2}}{a_{x}}+x}\right]$ (A.4)
Again, from equations (A.2) and (A.3) we can write as:
$\Bigg[\frac{c^{2}}{a_{x}}\left(1+\frac{a_{x}}{c^{2}}\,x\right)\Bigg]^{2}-c^{2}t^{2}$ $=\left(\frac{c^{2}}{a_{x}}\right)^{2}\Bigg\{1+\frac{a_{x}}{c^{2}}\,\bar{x}\Bigg\}^{2}$
$\times\Bigg[\cosh^{2}\left(\frac{a_{x}\,\bar{t}}{c}\right)-\sinh^{2}\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg]$
@@
$\text{or\,\,\,\,\,\,}\bar{x}$ $=\sqrt{\Bigg[\left(\frac{c^{2}}{a_{x}}+x\right)\Bigg]^{2}-c^{2}t^{2}}-\frac{c^{2}}{a_{x}}$ (A.5)
Equations (A.4) and (A.5) represents the inverse Rindler transformation between an accelerated frame $\bar{K}$ and a rest frame $K$ along the $X$-direction.
A.2 Inverse Rindler transformation for acceleration in arbitrary direction
@@
$\vec{r}_{\parallel}$ $=\frac{\left(\vec{r}\cdot\vec{a}\right)}{|\vec{a}|^{2}}\vec{a}\,\,\hskip 28.45274pt\text{and\,\,\,\,\,\,\,}\vec{r}_{\perp}=\vec{r}-\frac{\left(\vec{r}\cdot\vec{a}\right)}{|\vec{a}|^{2}}\vec{a}$ (A.7)
@@
$\vec{\bar{r}}_{\parallel}$ $=\Bigg[\sqrt{\left(\frac{\left(\vec{r}\cdot\vec{a}\right)}{|\vec{a}|^{2}}+\frac{c^{2}}{|\vec{a}|^{2}}\right)^{2}-\frac{c^{2}t^{2}}{|\vec{a}|^{2}}}-\frac{c^{2}}{|\vec{a}|^{2}}\Bigg]\vec{a}$ (A.9)
@@
$=\frac{c^{2}}{a}\tanh^{-1}\Bigg\{\frac{a\,t}{c\Big(\frac{\vec{a}}{c^{2}}\cdot\vec{r}+1\Big)}\Bigg\}$ (A.11)
and the spatial coordinate transformation can be obtained from (A.9) and (A.10) (by putting $|\vec{a}|=a$):
@@
$=\vec{r}-\Bigg[\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}-\Bigg\{\left(\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}+\frac{c^{2}}{a}\right)^{2}-c^{2}t^{2}\Bigg\}^{1/2}+\frac{c^{2}}{a}\Bigg]\frac{\vec{a}}{a}$ (A.12)
The equations (A.11) and (A.12) are the Rindler transformation of coordinates in arbitrary direction.
Now we consider that $\vec{a_{x}}$, $\vec{a_{y}}$ and $\vec{a_{z}}$ respectively are the $X$, $Y$ and $Z$ components of the acceleration $\vec{a}$ such that $a=\sqrt{a_{x}^{2}+a_{y}^{2}+a_{z}^{2}}$. Then we obtain from eqn. (A.11) as below:
$c\bar{t}$ $=\frac{c^{2}}{a}\tanh^{-1}\Bigg\{\frac{a\,t}{c\Big(\frac{a_{x}}{c^{2}}\,x+\frac{a_{y}}{c^{2}}\,y+\frac{a_{z}}{c^{2}}\,z+1\Big)}\Bigg\}$ (A.13)
and from (A.12):
$\vec{\bar{x}}+$ $\vec{\bar{y}}+\vec{\bar{z}}=\vec{x}+\vec{y}+\vec{z}-\Bigg[\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}-\Bigg\{\left(\frac{\left(\vec{r}\cdot\vec{a}\right)}{a}+\frac{c^{2}}{a}\right)^{2}-c^{2}t^{2}\Bigg\}^{1/2}+\frac{c^{2}}{a}\Bigg]\frac{\vec{a}}{a}$
@@
$-\frac{c^{2}t^{2}}{a^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg]\vec{a}_{z}$ (A.14)
@@
$+\Bigg\{\left(\frac{1}{a^{2}}\left(x\,a_{x}+y\,a_{y}+z\,a_{z}\right)+\frac{c^{2}}{a^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg]a_{x}$ (A.15)
@@
$+\Bigg\{\left(\frac{1}{a^{2}}\left(x\,a_{x}+y\,a_{y}+z\,a_{z}\right)+\frac{c^{2}}{a^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg]a_{y}$ (A.16)
@@
$+\Bigg\{\left(\frac{1}{a^{2}}\left(x\,a_{x}+y\,a_{y}+z\,a_{z}\right)+\frac{c^{2}}{a^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a^{2}}\Bigg]a_{z}$ (A.17)
@@
$c\bar{t}$ $=\frac{c^{2}}{a_{xy}}\tanh^{-1}\Bigg\{\frac{a_{xy}\,t}{c\Big(\frac{a_{x}}{c^{2}}\,x+\frac{a_{y}}{c^{2}}\,y+1\Big)}\Bigg\}$ (A.18)
$\bar{x}$ $=x+\Bigg[-\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\Bigg\{\left(\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}$
$\hskip 170.71652pt-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg]a_{x}$ (A.19)
$\bar{y}$ $=y+\Bigg[-\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\Bigg\{\left(\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}$
$\hskip 170.71652pt-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg]a_{y}$ (A.20)
@@
$c\bar{t}$ $=\frac{c^{2}}{a_{x}}\tanh^{-1}\Bigg\{\frac{c\,t}{\Big(\frac{c^{2}}{a_{x}}+x\Big)}\Bigg\}$ (A.22)
$\bar{x}$ $=\Bigg\{\left(x+\frac{c^{2}}{a_{x}}\right)^{2}-c^{2}t^{2}\Bigg\}^{1/2}-\frac{c^{2}}{a_{x}}$ (A.23)
$\bar{y}$ $=y$ (A.24)
$\text{and\,\,\,\,}\bar{z}$ $=z$ (A.25)
@@
A.3 Rindler transformation for the frames with non-zero initial velocity
@@
$ct$ $=\gamma\left(ct^{\prime}+\frac{v_{x}}{c}\cdot x^{\prime}\right),\hskip 14.22636ptx=\gamma\left(x^{\prime}+v_{x}\,t^{\prime}\right),\hskip 14.22636pty=y^{\prime}\,\,,\hskip 14.22636ptz=z^{\prime}$ (A.26)
@@
$ct^{\prime}$ $=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\text{,\,\,\,\,}x^{\prime}=\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}$ (A.27)
@@
Here it is to be noted that as stated above the accelerated frame $\bar{K}$ has zero initial velocity with respect to the frame $K^{\prime}$, however the frame $K^{\prime}$ itself is moving with uniform velocity $v_{x}$, therefore according to the observer in the rest frame $K$, the accelerated frame has non-zero initial velocity. The coordinate transformation relation between the inertial rest frame $K$ and the accelerated frame $\bar{K}$ can be obtained by substituting the expressions of the coordinates in equation (A.27) into equation (A.26) as follows:
$ct$ $=\gamma\Bigg[\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cdot\Bigg\{\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}\Bigg\}\Bigg]$ (A.28)
$x$ $=\gamma\Bigg[\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)-\frac{c^{2}}{a_{x}}+\frac{v_{x}}{c}\,\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg]$ (A.29)
$y$ $=\bar{y}$ (A.30)
$z$ $=\bar{z}$ (A.31)
@@
Equations (A.28) and (A.29) can be rewritten as:
$\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)$ $=\gamma\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\Bigg[\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg]$ (A.32)
$\text{and\,\,\,\,\,\,}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)$ $=\gamma\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)\Bigg[\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg]$ (A.33)
Now dividing (A.32) by (A.33) we obtain:
$\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}$ $=\frac{\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)}{\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)}=\frac{\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}}{1+\frac{v_{x}}{c}\cdot\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)}$ (A.34)
@@
$\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}=\frac{\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\tanh\left\{\tanh^{-1}\left(\frac{v_{x}}{c}\right)\right\}}{1+\tanh\left(\frac{a_{x}\,\bar{t}}{c}\right)\tanh\left\{\tanh^{-1}\left(\frac{v_{x}}{c}\right)\right\}}$
$\hskip 68.28644pt=\tanh\Bigg\{\frac{a_{x}\,\bar{t}}{c}+\tanh^{-1}\left(\frac{v_{x}}{c}\right)\Bigg\}$
or $\frac{a_{x}\,\bar{t}}{c}=\tanh^{-1}\Bigg\{\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}\Bigg\}-\tanh^{-1}\left(\frac{v_{x}}{c}\right)$ (A.35)
@@
$c\bar{t}=\frac{c^{2}}{a_{x}}\tanh^{-1}\left\{\frac{\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)-\frac{v_{x}}{c}\left(x+\gamma\frac{c^{2}}{a_{x}}\right)}{\left(x+\gamma\frac{c^{2}}{a_{x}}\right)-\frac{v_{x}}{c}\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)}\right\}$ (A.36)
Additionally, squaring both sides of equations (A.32) and (A.33) we can obtain as follows:
$\left(x+\gamma\frac{c^{2}}{a_{x}}\right)^{2}$ $-\left(ct+\gamma\frac{v_{x}c}{a_{x}}\right)^{2}=\gamma^{2}\left(\frac{c^{2}}{a_{x}}+\bar{x}\right)^{2}\Bigg[\Bigg\{\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)$
$+\frac{v_{x}}{c}\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg\}^{2}-\Bigg\{\sinh\left(\frac{a_{x}\,\bar{t}}{c}\right)+\frac{v_{x}}{c}\cosh\left(\frac{a_{x}\,\bar{t}}{c}\right)\Bigg\}^{2}\Bigg]$